OpenStudy (anonymous):

If Jim Roller can paint a room in 3 hours while John Brush would take 5 hours, how long will they take if they work together?

OpenStudy (anonymous):

i found this: http://answers.yahoo.com/question/index?qid=20070729140155AAUsPDt

OpenStudy (anonymous):

a;1/8 hour b;8/15 hour c; 7/8 hours

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

What class is this?

OpenStudy (anonymous):

math

OpenStudy (anonymous):

I know, but what level of math?

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

You need to make a system of equations.

OpenStudy (anonymous):

Let, \(t\) be the time it takes to paint the room. \(r\) is the rate that Roller to paints it, \(b\) is the rate that Brush paints it.

OpenStudy (anonymous):

We know that \[ 3r = 5b \]

OpenStudy (anonymous):

We also know that \[ (r+b)t = 3r = 5b \]

OpenStudy (anonymous):

First solve for \(r\) in terms of \(b\) \[ 3r=5b\implies r = \frac{5}{3}b \]Now plug it into equation 2 \[ \left(\frac{5}{3}b+b\right)t=5b \]

OpenStudy (anonymous):

Then solve for \(t\).

OpenStudy (raden):

shorcut way : 3*5/(3+5) = 15/8 long way :p v1+v2 = x/3 + x/5 = 8x/15 = x/(15/8) so, if they work together will take 15/8 hours

OpenStudy (raden):

no correct options there